OpenStudy (anonymous):

For what value of K will the following system of linear equations have a solution? Will give a medal!! 2x + 6y= 4 x + 5y= 6 kx +5y= -2

OpenStudy (anonymous):

First you figure the values of x and y from the the system of the first 2 equations. I will use elimination to solve for x and y through the first two given equations. 2x + 6y = 4 - 2x + 10y = 12 --------------- -4y = -8 y = 2 --> Now plug in this value in to one of the first equations and solve for x. 2x + 6(2) = 4 2x = 4 - 12 2x = -8 x = -4 Now we know that y = 2 and x = -4. Now we plug in these values in to the equation with k and solve for k. k(-4) + 5(2) = -2 -4k + 10 = -2 -4k = - 12 k = 3 Therefore, k = 3 satisfies the above system of equations.